A bass clarinet can be modeled as a 120-cm-long open-closed tube. A bass clarinet player starts playing in a 20∘C room, but soon the air inside the clarinet warms to where the speed of sound is 352 m/s.

What is the change in frequency? 
Answer #1

L = length of the pipe = 120 cm = 1.20 m

T = room temperature = 20

V = speed of sound at room temperature = 331.4 + 0.6 T = 331.4 + 0.6 x 20 = 343.4 m/s

frequency at T = 20 is given as

f = V/(4L) = 343.4 /(4 x 1.20) = 71.5 hz

at V' = 352 m/s

frequency is given as

f' = V'/(4L) = 352/(4 x 1.20) = 73.3 hz

Change in frequency = f' - f = 73.3 - 71.5 = 1.8 hz